# Pose Prior Learner: Unsupervised Categorical Prior Learning for Pose Estimation

### v2

@@ 6 Impact Statement
Appendix A4 PPL Results of Keypoint Detection on Occluded Images
Figure A1: PPL results of keypoint detection as a function of number of inference iterations on images with RandomMasking from Human3.6m. The “mask ratio” in the legend specifies the masked proportion on the $32\times 32=1024$ image patches.
Figure A2: PPL results of keypoint detection as a function of number of inference iterations on images with CenterMasking from Human3.6m. The “mask size” in the legend refers to the width and height of the masked region, on $1024$ image patches.
Appendix A6 Training Techniques
To ensure convergence and stability during the training of PPL, we introduce three gradient dettachment techniques: (1) To address the broken gradient issue during the quantization step from $G$ to $G^{\prime}$, we adopt the approach from VQ-VAE (Van Den Oord et al., 2017). Specifically, our PPL copies the gradients of $G^{\prime}$ to $G$ for backward propagation, allowing the gradients to flow through the quantization step. (2) The hierarchical memory $M$ is updated using an exponential moving average to smooth the gradient updates, particularly during the early stages of training when $G$ can be quite noisy. This approach helps stabilize the learning process and ensures that $M$ retains more reliable information over time. (3) For $M$ to learn effective representations of $G$, it requires an accurate estimation of $T^{\prime}$, which depends on a good prior $V$ distilled from $M$. This creates a chicken-and-egg problem that complicates training. To address this, we introduce two gradient detachments to separate the training processes. First, we detach the gradients from $T$ and train the keypoint transformation and image reconstruction pathway, as shown by the red arrows in Figure 2. Second, we detach the gradients from $T^{\prime}$ to train the memory encoder and decoder, $MIX_{\text{enc}}$ and $MIX_{\text{dec}}$, as indicated by the green arrows in Figure ALABEL:fig:mem_stru in Appendix A5.
Appendix A7 Additional Visualizations

### v3

@@ 5 Discussion
Ablations on loss components. Our method employs four loss components: Image Reconstruction Loss, Boundary Loss, Link Regularization Loss, and Keypoint Configuration Reconstruction Loss. The Image Reconstruction Loss provides the primary training signal, ensuring that the learned keypoints are semantically meaningful and spatially aligned with the object. The Keypoint Configuration Reconstruction Loss is critical for the successful training of the hierarchical memory. The Boundary Loss and the Link Regularization Loss are auxiliary constraints that mainly stabilize and accelerate convergence. Without the Boundary Loss, the network becomes unstable during early training. Replacing it with a sigmoid function causes a clear performance drop, increasing the normalized L2 error from 2.56 to 2.70 on Human3.6M at resolution 256×256. Similarly, removing the Link Regularization Loss leads to a measurable decline, with the error rising from 2.56 to 2.64. The comparison results are shown in Table A3.
Table A3: Ablation studies on Human3.6M at resolution $256\times 256$. Numbers indicate normalized L2 error (lower is better). Best is in bold.
Ablation Results Normalized L2 Error
PPL-full (ours) 2.56
w/o Boundary Loss training unstable
Boundary Loss $\to$ Sigmoid 2.70
w/o Link Regularization Loss 2.64
PPL-1MemBank 2.62
Removal of hierarchy in the memory. We introduce an ablated variant of PPL using only one memory bank, referred to as PPL-1MemBank. Below, we introduce its implementation. If each pose’s keypoints were encoded into 34 vectors sharing the same embedding space, mean pooling these embeddings would not effectively distill the keypoint prior. Instead, PPL-1MemBank encodes all keypoints of a pose into a single vector, ensuring that all vectors share a common embedding space. To maintain comparable capacity, PPL-1MemBank uses a single memory bank with 512 vectors, equivalent to PPL’s 34 memory banks with 16 vectors each. The keypoint prior is then obtained by mean pooling across this single memory bank. We evaluated PPL-1MemBank against PPL, with results summarized in the last row of Table A3. On the Human3.6m dataset, PPL-1MemBank achieves a normalized L2 error of 2.72, compared to 2.56 with PPL’s multi-bank approach. This indicates that multiple smaller memory banks improve pose estimation performance.
PPL trained with randomly masked frames as reference images on video datasets. We also train PPL on Human3.6m and Taichi by treating video frames as independent images and using randomly masked images as $I_{ref}$. We denote this setting as PPL*. The results are reported in Table A4. Notably, even with masked images as $I_{ref}$, PPL* consistently outperforms AutoLink. This highlights that the effectiveness of PPL stems primarily from its pose prior, rather than from cross-frame reconstruction.
Table A4: Results of Training PPL* using the randomly selected and masked video frames as $I_{ref}$ on Human3.6M and Taichi video datasets. We use resolution 256×256 for Taichi and 128×128 for Human3.6m. We use the evaluation metrics as reported in Table 1. Best is in bold.
Methods Human3.6m Taichi
AutoLink 2.76 316.10
PPL (ours) 1.92 293.35
PPL* (ours) 2.23 298.60